Icy Candy announces a 1 for 8 bonus issues. Icing Candy shares are trading at $9.00 before the bonus issue.
a. Calculate the theoretical price of Icing Candy’s shares immediately after the bonus issue.
b. Casper has 1,000 shares in Icy Candy before Icing Candy announced the 1 for 8 bonus issue.
i. How many bonus shares will Casper entitle to?
ii. Find the value of Casper’s stockholding in Icy Candy before and after the bonus issue.
1
Expert's answer
2019-12-09T09:27:17-0500

a. The theoretical price of Icing Candy’s shares immediately after the bonus issue is: 9×8/(8 + 1) = $8.

b. If Casper has 1,000 shares in Icy Candy before Icing Candy announced the 1 for 8 bonus issue, then:

i. Casper will get 1000/8 = 125 bonus shares.

ii. Find the value of Casper’s stockholding in Icy Candy before the bonus issue is:

1000×9 = $9,000.

After the bonus issue the value of Casper’s stockholding in Icy Candy is:

(1,000 + 125)×8 = $9,000.